OpenStudy (anonymous):

two planes alpha: 2x+3y-2z-3=0 and beta: 5x-3y+3z-5=0 intersect each other on the line t. with three planes i would be able to find the answer but with only two planes?

OpenStudy (amistre64):

three planes make a point; it how you solve for the coeefs in a parabola..

OpenStudy (amistre64):

simply put z over to the other side and treat it as a constant

OpenStudy (amistre64):

2x+3y = 2z +3 5x -3y =-3z+5

OpenStudy (amistre64):

2x+3y = 2z +3 5x -3y =-3z+5 -------------- 7x = -z+8 x = (-z+8)/7 ; now solve for y :)

OpenStudy (amistre64):

the line formed will be defined by a scaled vector and a point

OpenStudy (anonymous):

so i do need a third plane for a clear point?

OpenStudy (amistre64):

for a "POINT" yes; but to find the "line" that they meet at? no

OpenStudy (anonymous):

thank you very much.

OpenStudy (amistre64):

yep, youre welcome :)